# London, UK Dual Branch VideoMamba for Violence Detection

### v1

@@ Abstract
The rapid proliferation of surveillance cameras has increased the demand for automated violence detection. While CNNs and Transformers have shown success in extracting spatio-temporal features, they struggle with long-term dependencies and computational efficiency. We propose Dual Branch VideoMamba with Gated Class Token Fusion (GCTF), an efficient architecture combining a dual-branch design and a state-space model (SSM) backbone where one branch captures spatial features, while the other focuses on temporal dynamics, with continuous fusion via a gating mechanism. We also present a new benchmark by merging RWF-2000, RLVS, and VioPeru datasets in video violence detection, ensuring strict separation between training and testing sets. Our model achieves state-of-the-art performance on this benchmark offering an optimal balance between accuracy and computational efficiency, demonstrating the promise of SSMs for scalable, real-time surveillance violence detection.
@@ 1 Introduction
Figure 1: (a) Accuracy vs FLOPS, (b) Accuracy vs Parameters.
The increased installation of surveillance cameras in public and private spaces driven by advancements in low-cost imaging technology has led to a dramatic increase in the amount of surveillance video data generated daily [Ullah et al.(2023)Ullah, Obaidat, Ullah, Muhammad, Hijji, and Baik, Omarov et al.(2022)Omarov, Narynov, Zhumanov, Gumar, and Khassanova]. Since continuous human monitoring of these numerous surveillance video feeds is neither practical nor reliable to detect violent behaviors, there is a significant interest in developing deep learning models that can reliably and efficiently detect violence in surveillance videos.
Recent advancements in computer vision have leveraged deep convolutional neural networks (CNNs) and, more recently, Transformer-based architectures to extract rich spatio-temporal features from video data [Ji et al.(2012)Ji, Xu, Yang, and Yu, Arnab et al.(2021)Arnab, Dehghani, Heigold, Sun, Lučić, and Schmid]. However, while 3D CNNs excel at capturing local patterns, they often struggle with the long-term dependencies that are crucial for understanding extended video sequences. On the other hand, Transformer models, although powerful in modeling global context, suffer from quadratic computational complexity, which limits their scalability [Arnab et al.(2021)Arnab, Dehghani, Heigold, Sun, Lučić, and Schmid].
In this work, we propose the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) model which is designed to address these limitations with the help of an efficient state-space model (SSM) backbone. Based on the recent VideoMamba framework [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao], we introduce a dual-stream processing architecture that continuously fuses the learned information from both branches through a gated class token fusion mechanism between the two branches. Our model processes two parallel streams: one dedicated to capturing fine-grained spatial cues and the other focused on temporal information while fusing learned features of each branch through fusing the CLS tokens pertaining to each branch continuously with the help of a gated learnable parameter. This fusion allows the network to effectively detect violent activities even in scenarios where subtle motion differences are critical.
In addition to the novel architectural design, we address a key challenge in surveillance violence detection dataset diversity and generalizability by introducing an amalgamated dataset that combines three established surveillance benchmarks: Real World Fighting (RWF-2000) [Cheng et al.(2021)Cheng, Cai, and Li], Real Life Violence Situations (RLVS) [Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab], and VioPeru [Huillcen Baca et al.(2024)Huillcen Baca, Palomino Valdivia, and Gutierrez Caceres]. This integrated surveillance dataset provides a broader range of violent scenarios, improving the generalizability of our approach across different environments and recording conditions. Our contributions are threefold:
@@
    We propose a dual-branch VideoMamba architecture that leverages the efficiency of state-space models while explicitly decoupling spatial and temporal reasoning through two specialized pipelines with different scanning strategies. Unlike prior dual-encoder SSM models, our architecture is tailored for classification tasks without memory modules, and enables cross-branch interaction at the semantic (class token) level.
@@
    We introduce a novel continuous Gated Class Token Fusion (GCTF) mechanism that enables adaptive integration of the class tokens from the dual streams, ensuring a balanced and adaptive fusion of learned spatial and temporal features.
@@
    We curate an integrated surveillance violence detection dataset by amalgamating RWF-2000, RLVS, and VioPeru, while rigorously preventing data leakage between training and testing sets where extensive experiments demonstrate that our Dual Branch VideoMamba with GCTF model achieves state-of-the-art performance in surveillance violence detection with respect to the combined dataset while increasing computational efficiency as evident in Fig. 1 and Table 2, making it a promising solution for real-world surveillance applications.
@@ 2 Related Work
Research in automated violence detection has primarily followed two paradigms: anomaly detection and action recognition. In the anomaly detection paradigm [Pu et al.(2024)Pu, Wu, Yang, and Wang, Chen et al.(2023)Chen, Liu, Zhang, Fok, Qi, and Wu], violent events are modeled as rare deviations from normal activity. While, this approach has shown promise in controlled environments, they have often failed to capture the complex context in which violence occurs in real-world [Sultani et al.(2018)Sultani, Chen, and Shah, Chen et al.(2023)Chen, Liu, Zhang, Fok, Qi, and Wu, Tian et al.(2021)Tian, Pang, Chen, Singh, Verjans, and Carneiro].
@@
In contrast, the action recognition paradigm treats violence detection as a supervised classification problem where the focus is on learning discriminative features for violent versus non-violent actions. Early methods utilized 3D CNNs to capture spatio-temporal features directly from video clips [Ullah et al.(2019)Ullah, Ullah, Muhammad, Haq, and Baik]. Subsequent research [Su et al.(2020)Su, Lin, Zhu, and Wu] introduced techniques such as skeleton-based action recognition, where human pose estimation is used to generate 3D skeleton representations, and Graph CNNs are employed to model the interactions between individuals. [Su et al.(2020)Su, Lin, Zhu, and Wu] is one of the first papers to evaluate performance on a real-world surveillance violence detection data set (RWF-2000) [Cheng et al.(2021)Cheng, Cai, and Li] where almost all the previous literature were evaluated on non-surveillance based datasets such as the Hockey Fight dataset [Bermejo Nievas et al.(2011)Bermejo Nievas, Deniz Suarez, Bueno García, and Sukthankar]. [Garcia-Cobo and SanMiguel(2023), Asad et al.(2022)Asad, Jiang, Yang, Tu, and Malik, Islam et al.(2021)Islam, Rukonuzzaman, Ahmed, Kabir, and Farazi] have used deep architectures comprising of two simultaneous pipelines to extract different types of features similarly to [Feichtenhofer et al.(2019)Feichtenhofer, Fan, Malik, and He] in action recognition space. These methods have shown robust performance on benchmark datasets, but still face challenges in effectively modeling long-term dependencies due to their reliance on standard convolutional operations.
@@
Lately, Transformer-based models have been applied to violence detection by leveraging self-attention mechanisms to capture global context [Li et al.(2023a)Li, Yang, and Liang]. Recently, models such as [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh] have incorporated architectures that amalgamate convolution and self-attention mechanisms to balance computational efficiency with representational power to effectively capture local and global dependencies in the context of violence detection inspired by video action recognition models such as [Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao]. Although such models have achieved impressive performance, their quadratic complexity in attention operations can hamper their scalability [Li et al.(2023c)Li, Wang, Zhang, Gao, Song, Liu, Li, and Qiao, Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao].
@@
The emergence of efficient subquadratic-time architectures with state-space models (SSMs) has opened a new avenue for addressing these challenges [Gu and Dao(2024), Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. VideoMamba [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] extends the work of [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] in the video recognition space where it replaces the traditional self-attention mechanism with a linear-complexity state-space module, enabling efficient processing of video sequences without sacrificing the ability to capture long-range dependencies. Similarly, [Li et al.(2024b)Li, Zhao, Yang, Liu, Sheng, Zeng, Wang, Wu, and Jiang] has introduced another SSM architecture for video anomaly detection by effectively modeling spatial-temporal normality through a Spatial-Temporal interaction module. Our approach mitigates the drawbacks of traditional CNNs and Transformers while benefiting from the computational efficiency of SSMs. To our knowledge, this paper is the first to investigate the use of SSMs for video violence detection.
@@ 3 Proposed Method
Our design builds upon the VideoMamba framework and its efficient state-space model modules. We employ two parallel pipelines—each implementing a variant of VideoMamba with distinct scanning strategies—with the intention of separately extracting spatial and temporal features from video inputs, combined with a cropping module to focus on human-interactions. We introduce a novel Gated Class Token Fusion (GCTF) mechanism that combines information between the two branches. GCTF is performed at each layer in the network, providing a form of continuous fusion. The enriched feature representation is combined from the Spatial-First Scanning Branch to the Temporal-First Scanning Branch to form a unified representation for final classification.
@@ 3.1 Overview of Mamba Architectures
The Mamba family of architectures rethinks the standard self-attention mechanism by leveraging efficient state-space models (SSMs) for long-range dependency modeling  [Gu and Dao(2024)]. The key idea is to view a sequence as the output of a continuous dynamical system, where the evolution of a hidden state is governed by ordinary differential equations [Gu and Dao(2024)]. In its continuous form, a state-space model is given by:
@@
where $x(t)$ is the input at time $t$, $h(t)\in\mathbb{R}^{N}$ is the hidden state, and where $\mathbf{A}\in\mathbb{R}^{N\times N}$ represents the evolutionary matrix of the system and $\mathbf{B}\in\mathbb{R}^{N\times D}$, and $\mathbf{C}\in\mathbb{R}^{d\times N}$ are projection matrices. To process discrete token sequences, this continuous system is approximated via discretization (commonly using a zero-order hold), which includes a timescale parameter $𝚫{\mathbf{\Updelta}}$ to transform the continuous parameters ${\mathbf{A}},{\mathbf{B}}$ to their discrete learnable counterparts $\overline{{\mathbf{A}}},\overline{{\mathbf{B}}}$:
@@ 3.2 Dual Branch VideoMamba with GCTF Architecture
We introduce our novel architecture, the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) for Violence Detection in surveillance videos as shown in Fig. 2. The architecture contains four main components, namely: (a) Cropping Module; (b) Branch-1 (Spatial-First Scanning); (c) Branch-2 (Temporal-First Scanning); and (d) Final Fusion Block; inspired by the motivational factors discussed in the preceding section.
Figure 2: Part (a): The overall Dual Branch VideoMamba with GCTF Architecture with: 1) Cropping module to detect people and crop spatially the area of focus; 2) Branch-1 and Branch-2 which implement the VideoMamba architecture in parallel with lateral connections for continuous Gated Class Token Fusion (GCTF); 3) Final Fusion Block. Part (b): Bi-Directional Vision Mamba Encoder Block [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang].
@@ 3.2.1 Cropping Module
To focus on human actions, our cropping mechanism extracts the region that encompasses all detected people in each frame to guide the network, based on the observation that violent incidents typically involve multiple individuals. More details on this are discussed in Section 6.1 of supplementary material. This approach, also adopted in CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh], enhances the model’s ability to accurately identify violent behavior as evident in the ablation study.
@@ 3.2.2 Dual-Branch Architecture Overview
Figure 3: (a) Spatial-First Scanning (b) Temporal-First Scanning.
@@
This branch reorganizes the tokens such that the spatial layout within each frame is prioritized. In other words, tokens are grouped and ordered by their spatial coordinates before being concatenated across frames as depicted in Fig. 3 (a). This strategy is aimed at learning fine-grained local spatial features that are critical for recognizing visual cues. This Spatial-First scan in this branch is performed bi-directionally.
@@
Here, the tokens are reorganized based on the frame and then stacked along the spatial dimension maintaining their natural temporal order so that the sequential progression of frames is preserved as depicted in Fig. 3 (b). This ordering can facilitate the model to capture dynamic motion patterns and global temporal dependencies pertaining to each of those tokens, which are essential for detecting violent actions and changes over time. Similar to the Branch-1, this scan too is performed bi-directionally.
This strategy enables leveraging the pre-trained weights of a vanilla VideoMamba to initialize each branch in our architecture. Pre-trained initialization is beneficial because it provides a robust starting point, accelerates convergence during training, and enhances overall performance by transferring learned representations from extensive prior training  [Li et al.(2023c)Li, Wang, Zhang, Gao, Song, Liu, Li, and Qiao]. During the forward pass, Branch-1 processes its input video to produce a sequence of features, including a dedicated class (CLS1) token for each layer. Meanwhile, Branch-2 receives two inputs: the video embedding tokens padded with the CLS2 token and additional intermediate (CLS1) token features from each block of Branch-1 (excluding the final block). This second input allows Branch-2 to incorporate spatial context learned from Branch-1 into its feature learning to produce better results as evident by the results in Table 5 of the ablation study.
@@ 3.2.3 Gated Class Token Fusion (GCTF) between Branch-1 and Branch-2
$\text{CLS}_{\text{fused(l)}}^{2}=\sigma^{\prime}_{l}\odot\text{CLS}_{l}^{2}+\left(1-\sigma^{\prime}_{l}\right)\odot\text{CLS}_{l}^{1},$ (7)
where $\sigma_{l}\in\mathbb{R}^{d}$ is a learnable parameter vector for block $l$, and it is passed through the Sigmoid function to ensure that the gate value lie between 0 and 1 before performing element wise multiplication denoted by $\odot$ with CLS tokens from each branch to dynamically weigh the best contributions. Unlike typical fusion schemes that operate at a single depth or combine all tokens, GCTF selectively merges class-level semantics across branches at every depth, enabling layerwise refinement and preventing early overcommitment to either modality.

### v2

# Dual Branch VideoMamba with Gated Class Token Fusion for Violence Detection

@@ Abstract
The rapid proliferation of surveillance cameras has increased the demand for automated violence detection. While CNNs and Transformers have shown success in extracting spatio-temporal features, they struggle with long-term dependencies and computational efficiency. We propose Dual Branch VideoMamba with Gated Class Token Fusion (GCTF), an efficient architecture combining a dual-branch design and a state-space model (SSM) backbone where one branch captures spatial features, while the other focuses on temporal dynamics. The model performs continuous fusion via a gating mechanism between the branches to enhance the model’s ability to detect violent activities even in challenging surveillance scenarios. We also present a new benchmark by merging RWF-2000, RLVS, SURV and VioPeru datasets in video violence detection, ensuring strict separation between training and testing sets. Experimental results demonstrate that our model achieves state-of-the-art performance on this benchmark and also on DVD dataset which is another novel dataset on video violence detection, offering an optimal balance between accuracy and computational efficiency, demonstrating the promise of SSMs for scalable, near real-time surveillance violence detection.
@@ 1 Introduction
The increased installation of surveillance cameras in public and private spaces driven by advancements in low-cost imaging technology has led to a dramatic increase in the amount of surveillance video data generated daily [44, 30]. Since continuous human monitoring of these numerous surveillance video feeds is neither practical nor reliable to detect violent behaviors, there is a significant interest in developing deep learning models that can reliably and efficiently detect violence in surveillance videos.
Figure 1: Model comparison on the combined dataset showing Top-1 Accuracy vs. FLOPS, with marker size proportional to the number of parameters.
Recent advancements in computer vision have leveraged deep convolutional neural networks (CNNs) and, more recently, Transformer-based architectures to extract rich spatio-temporal features from video data [18, 3]. However, while 3D CNNs excel at capturing local patterns, they often struggle with the long-term dependencies that are crucial for understanding extended video sequences. On the other hand, Transformer models, although powerful in modeling global context, suffer from quadratic computational complexity, which limits their scalability [3].
In this work, we propose the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) model which is designed to address these limitations with the help of an efficient state-space model (SSM) backbone. Based on the recent VideoMamba framework [26], we introduce a dual-stream processing architecture that continuously fuses the learned information from both branches through a gated class token fusion mechanism between the two branches. Our model processes two parallel streams: one dedicated to capturing fine-grained spatial cues and the other focused on temporal information while fusing learned features of each branch through fusing the CLS tokens pertaining to each branch continuously with the help of a gated learnable parameter. This fusion allows the network to effectively detect violent activities even in scenarios where subtle motion differences are critical.
In addition to the novel architectural design, we address a key challenge in surveillance violence detection dataset diversity and generalizability by introducing an amalgamated dataset that combines four established open-source strongly labeled video violence benchmarks: Real World Fighting (RWF-2000) [9], Real Life Violence Situations (RLVS) [36], Vision-based Fight Detection From Surveillance Cameras (SURV) [2] and VioPeru [16]. This integrated surveillance dataset provides a broader range of violent scenarios, improving the generalizability of our approach across different environments and recording conditions. Our contributions are threefold:
@@
    Dual-Branch State-Space Architecture: We introduce the first state-space-based design for violence detection. Our dual-branch architecture explicitly decouples spatial and temporal reasoning while enabling continuous interaction at the semantic level through class tokens. This differs from prior dual-encoder or CNN+Transformer designs, which rely on late fusion or heavy attention mechanisms.
@@
    Gated Class Token Fusion (GCTF): We propose a novel continuous fusion mechanism that adaptively integrates class tokens across network layers. This enables layer-wise refinement of spatial and temporal representations while avoiding premature overcommitment to either branch.
@@
    We curate an integrated surveillance violence detection dataset by amalgamating RWF-2000, RLVS, SURV and VioPeru, while rigorously preventing data leakage between training and testing sets where extensive experiments demonstrate that our Dual Branch VideoMamba with GCTF model achieves state-of-the-art performance in surveillance violence detection with respect to the combined dataset and the novel DVD dataset [22] while increasing computational efficiency as evident in Fig. 1 and Tab. 1, making it a promising solution for real-world surveillance applications.
@@ 2 Related Work
Research in automated violence detection has primarily followed two paradigms: anomaly detection and action recognition. In the anomaly detection paradigm [33, 8], violent events are modeled as rare deviations from normal activity. While this approach has shown promise in controlled environments, methods often fail to capture the complex context in which violence occurs in real-world [39, 8, 41].
@@
In contrast, the action recognition paradigm treats violence detection as a supervised classification problem where the focus is on learning discriminative features for violent versus non-violent actions. Early methods utilized 3D CNNs to capture spatio-temporal features directly from video clips [43, 37]. Subsequent research [38] introduced techniques such as skeleton-based action recognition, where human pose estimation is used to generate 3D skeleton representations, and Graph CNNs are employed to model the interactions between individuals.  [38] is one of the first papers to evaluate performance on a real-world surveillance violence detection dataset (RWF-2000) [9] where almost all the previous literature were evaluated on non-surveillance based datasets such as the Hockey Fight dataset [5].  [13, 4, 17, 46] employed deep architectures comprising of two simultaneous pipelines to extract different types of features similarly to [11] in action recognition space. These methods have shown robust performance on benchmark datasets, but still face challenges in effectively modeling long-term dependencies due to their reliance on standard convolutional operations.
@@
Lately, Transformer-based models have been applied to violence detection by leveraging self-attention mechanisms to capture global context [23]. Models such as [35] have incorporated architectures that amalgamate convolution and self-attention mechanisms to balance computational efficiency with representational power to effectively capture local and global dependencies in the context of violence detection inspired by video action recognition models such as [25, 24]. Although such models have achieved impressive performance, their quadratic complexity in attention operations can hamper their scalability [24, 25].
@@
The emergence of efficient subquadratic-time architectures with state-space models (SSMs) has opened a new avenue for addressing these challenges [14, 48]. VideoMamba [26] extends the work of [48] in the video recognition space where it replaces the traditional self-attention mechanism with a linear-complexity state-space module, enabling efficient processing of video sequences without sacrificing the ability to capture long-range dependencies. Similarly, [27] introduced another SSM architecture for video anomaly detection by effectively modeling spatial-temporal normality through a Spatial-Temporal interaction module. This approach mitigates the drawbacks of traditional CNNs and Transformers while benefiting from the computational efficiency of SSMs. To our knowledge, this paper is the first to investigate the use of SSMs for video violence detection.
@@ 3 Proposed Method
Our design builds upon the VideoMamba framework and its efficient state-space model modules. We employ two parallel pipelines, each implementing a variant of VideoMamba with distinct scanning strategies, with the intention of separately extracting spatial and temporal features from video inputs, combined with a cropping module to focus on human-interactions. We introduce a novel Gated Class Token Fusion (GCTF) mechanism that combines information between the two branches. GCTF is performed at each layer in the network, providing a form of continuous fusion. The enriched feature representation is combined from the Spatial-First Scanning Branch to the Temporal-First Scanning Branch to form a unified representation for final classification.
@@ 3.1 Overview of Mamba Architectures
The Mamba family of architectures rethinks the standard self-attention mechanism by leveraging efficient state-space models (SSMs) for long-range dependency modeling [14]. The key idea is to view a sequence as the output of a continuous dynamical system, where the evolution of a hidden state is governed by ordinary differential equations [14]. In its continuous form, a state-space model is given by:
@@
[…]
@@
Vanilla VideoMamba:
The original VideoMamba model [26] applies the state-space formulation discussed above from Vision Mamba [48] to video sequences. In this architecture, the same bi-directional Mamba block used in Vision Mamba [48] depicted in Fig. 2 (b) has been used to process spatio-temporal information of the video sequence. While Vision Mamba adapts the approach to visual data by incorporating bi-directional processing and explicit positional embeddings, VideoMamba focuses on efficiently modeling spatio-temporal dynamics by adding an extra temporal embedding. These innovations form the foundation for our proposed dual-branch approach, which further exploits the complementary strengths of spatial and temporal feature learning.
@@ 3.2 Dual Branch VideoMamba with GCTF Architecture
We introduce a novel architecture, the Dual Branch VideoMamba with Gated Class Token Fusion (GCTF) for Violence Detection in surveillance videos as shown in Fig. 2. The architecture contains four main components, namely: (a) Cropping Module; (b) Branch-1 (Spatial-First Scanning); (c) Branch-2 (Temporal-First Scanning); and (d) Final Fusion Block; inspired by the motivational factors discussed in the preceding section.
@@ 3.2.1 Cropping Module
To focus on human actions, the cropping mechanism extracts the region that encompasses all detected people in each frame, based on the observation that violent incidents typically involve multiple individuals. As seen in Fig. 2 (a), by computing the maximum bounding box around all people, the network is guided to the most informative spatial areas while retaining the original frame if no individuals are detected. This approach, also adopted in CUE-Net [35], enhances the model’s ability to accurately identify violent behavior as evident in the ablation study. More details on this are discussed in Sec. 5.1 of supplementary material.
Figure 2: Part (a): The overall Dual Branch VideoMamba with GCTF Architecture with: 1) Cropping module to detect people and crop spatially the area of focus; 2) Branch-1 and Branch-2 which implements VideoMamba architecture in two parallel pathways with lateral connections between parallel layers implementing continuous Gated Class Token Fusion (GCTF); 3) Final Fusion Block. Part (b): Bi-Directional Vision Mamba Encoder Block [48] used in the above architecture as well as in VideoMamba architecture
@@ 3.2.2 Dual-Branch Architecture Overview
This branch reorganizes the tokens such that the spatial layout within each frame is prioritized. Tokens are grouped and ordered by their spatial coordinates before being concatenated across frames as depicted in Fig. 3 (a). This strategy is aimed at learning fine-grained local spatial features that are critical for recognizing visual cues. This Spatial-First scan in this branch is performed bi-directionally.
@@
Here, the tokens are reorganized based on the frame and then stacked along the spatial dimension maintaining their natural temporal order so that the sequential progression of frames is preserved as depicted in Fig. 3 (b). This ordering can facilitate the model to capture dynamic motion patterns and global temporal dependencies pertaining to each of those tokens, which are essential for detecting violent actions and changes over time. Similar to Branch-1, this scan too is performed bi-directionally. This strategy enables leveraging the pre-trained weights of a vanilla VideoMamba to initialize each branch in our architecture. Pre-trained initialization is beneficial because it provides a robust starting point, accelerates convergence during training, and enhances overall performance by transferring learned representations from extensive prior training [24].
During the forward pass, Branch-1 processes its input video to produce a sequence of features, including a dedicated class (CLS1) token for each layer. Meanwhile, Branch-2 receives two inputs: the video embedding tokens padded with the CLS2 token and additional intermediate (CLS1) token features from each block of Branch-1 (excluding the final block). This second input allows Branch-2 to incorporate spatial context learned from Branch-1 into its feature learning to produce better results as evident by the results in Tab. 5 of the ablation study. To enhance information flow, residual skip connections are also incorporated among the blocks within each branch.
@@ 3.2.3 Gated Class Token Fusion (GCTF) between Branch-1 and Branch-2
$\text{CLS}_{\text{fused(}l\text{)}}^{2}=\sigma^{\prime}_{l}\odot\text{CLS}_{l}^{2}+\left(1-\sigma^{\prime}_{l}\right)\odot\text{CLS}_{l}^{1},$ (7)
where $\sigma_{l}\in\mathbb{R}^{d}$ is a learnable parameter vector for block $l$, and it is passed through the Sigmoid function to ensure that the gate value lies between 0 and 1 before performing element-wise multiplication denoted by $\odot$ with CLS tokens from each branch to dynamically weigh the best contributions. Unlike typical fusion schemes that operate at a single depth or combine all tokens, GCTF selectively merges class-level semantics across branches at every depth, enabling layer wise refinement and preventing early overcommitment to either modality.